The Test score for a class are normally distributed . what is the the probability a student scored above 90


1
Expert's answer
2022-05-17T08:12:46-0400

Given "\mu=75, \sigma=10."


"P(X>90)=1-P(Z\le\dfrac{90-75}{10})"

"=1-P(Z\le1.5)=0.0668"